Large capacitors can hold a potentially dangerous charge long after a circuit has been turned off, so it is important to make sure they are discharged before you touch them. Suppose a 120 μFcapacitor from a camera flash unit retains a voltage of 180V when an unwary student removes it from the camera.

If the student accidentally touches the two terminals with his hands, and if the resistance of his body between his hands is 1.8 kΩ, for how long will the current across his chest exceed the danger level of 50 mA?

Express your answer with the appropriate units.
Answer #1

Current, I = (V/R) e^-(t/RC)

0.050 = (V/R) e^-(t/RC)

0.05 x 1800/180 = e^-(t/RC)

e^-(t/RC) = 0.5

-t/RC = - 0.693

t = 0.693 x 1800 x 120 x 10^-6 = 149.7 ms

Comment in case any doubt please rate my answer....